# Testing and Enhancing Multi-Agent Systems for Robust Code Generation

### v1

@@ 6.2. Overview of Our Repairing Method
As illustrated in Fig. 2, upon receiving an input requirement, we first mutate the requirement and generate multiple semantic-equivalent expressions to reduce the chance that MASs misunderstand specific expressions. For each generation, the planner generates the original plan and forwards it to the monitor, a new agent we introduce to mitigate the planner-coder gap. After receiving the plan from the planner, the monitor agent interprets the plan by providing detailed explanations for core concepts, edge cases, complex logic, relational phrases, and conditional judgments (see Sec. 5), which compensate for the information loss. Subsequently, the coder implements code based on this interpreted plan. In addition, the monitor also checks the implementation to enhance the alignment with the interpreted plan and requests regeneration when misalignments are detected.
@@ 6.4. Monitor Agent Insertion
To enhance plan quality and strengthen planner-coder alignment, we introduce a monitor agent between planner and coder. This new agent in the whole MAS pipeline performs two critical tasks: plan interpretation and code check. The monitor agent serves as a crucial mechanism to compensate for the information loss inherent in the process of code generation.
@@
To effectively address this issue and enhance plan quality, we prompt the inserted monitor agent to interpret the plan generated by the planner to make it more comprehensible to the coder. The monitor agent aims to enhance the clarity and completeness of the planned coding logic and mitigate the planner-coder gap, rather than correcting the semantic content of the plan. Specifically, the monitor interprets the decomposed requirement $\mathcal{R}$ and coding logic $\mathcal{L}$ with detailed explanations and clarifications targeting the five error patterns identified in Sec. 5, producing an interpreted plan $p^{\prime}$ that bridges the semantic gap between planning and coding. Fig. 3 presents an illustrative example of our prompt for the monitor agent. By interpreting and expanding the plan $p$ into $p^{\prime}$, the monitor effectively reduces the information loss between the coding logic and final implementation, which ensures that critical implementation details lost in Eq. 34 are restored before code generation. We employ few-shot prompting with carefully crafted examples to guide the interpretation process, focusing on the fundamental components of programming plans—such as core concepts, logic flows, and edge cases—rather than being tailored to any specific problem domain. This task-agnostic design provides general applicability across different code generation scenarios.
Clarification. Note that one might consider incorporating the monitor’s interpretation logic directly into the planner’s prompt. However, our proposed approach of a separate monitor agent offers advantages in terms of modularity and scalability. Rather than modifying the planner’s core functionality, this separation of concerns allows the monitor to be seamlessly integrated as a “quality gate” into various existing MAS architectures without altering their fundamental agent logic or task distribution. This design promotes reusability, easier maintenance, and addresses robustness issues across different MAS implementations without requiring architecture-specific modifications. We also follow the practice of consistent backend in a MAS to build the monitor agent with the backend LLMs of the MAS. The complete prompt is available in our artifact. [^3] https://github.com/anonymous-for-submit/MAS_testing_repairing/tree/main/More_results/Monitor
@@
Instead of using the testcase-based testing approaches that require dynamic execution with proper running environment, we directly prompt LLMs to perform static inspection and evaluate code compliance with the detailed specifications in $p^{\prime}$, which has been shown to be effective with the rich contextual information provided by the interpreted plan (Wang et al., 2024b, a). After the coder produces the solution code, we feed both the interpreted plan and the code back to the monitor to check whether the code complies with the plan regarding the different aspects of interpretation, as shown in Fig. 3(c). If the monitor identifies mismatches between code implementation and the plan, it returns the implementation to the coding agent for revision. By checking alignment between the interpreted plan $p^{\prime}$ and generated code $c$, the monitor creates a validation loop that detects and corrects information drift, effectively reducing the information loss that would otherwise accumulate through unchecked transformations. For efficiency, we use zero-shot prompting for code check and limit the process to execute only once.
@@ 7.1. RQ1: Can our repairing method effectively mitigate failures identified in fuzzing?
MAS Backend LLM HumanEval ET MBPP ET CodeContest CoderEval
Total Solved Ratio Total Solved Ratio Total Solved Ratio Total Solved Ratio
@@
Category EP-1 EP-2 EP-3 EP-4 EP-5 Overall Plan Logic Error Invalid
Solved (%) 80.0 83.3 80.0 100.0 85.7 83.9 66.7 0.0
@@
Dataset characteristics also affect repair performance. HumanEval and CoderEval generally achieves better repair performance, while CodeContest exhibits relatively lower repair rates. This disparity reflects the inherent question length differences: MASs generate concrete, focused plans that are straightforward to interpret for HumanEval’s concise questions, whereas CodeContest’s complex problems lead to verbose plans that challenge effective interpretation, which may exacerbate the planner-coder gap. A potential improvement could involve summarizing verbose questions to avoid excessively long questions or plans.
To further understand the strengths of our method, Table 5 presents the repair results for different error patterns identified in SCCG (GPT-3.5). Our repairing method demonstrates remarkable effectiveness in handling failures originating from the planner-coder gap, successfully resolving 83.9% of such cases. Among the five EPs illustrated in Sec. 5, our repairing method performs best on failures originating from relational phrases, suggesting that emphasizing such phrases can effectively enhance the coder’s understanding of these expressions. As expected, our method shows limited effectiveness on plan logic errors, since these represent the reasoning flaws of the planner rather than communication issues that our interpretation mechanism is designed to address.
@@ 7.2. RQ2: How do different components contribute to the overall performance?
We conduct ablation studies to investigate the effectiveness of multi-prompt generation and monitor agent insertion in our repairing method. For monitor agent insertion, we separately examine “Plan Interpretation” and “Code Check” to provide in-depth analysis. We present results with GPT-3.5 as the backend LLM, as it represents a typical configuration with other LLMs exhibiting similar patterns. Complete results are available in our artifact.  [^4] https://github.com/anonymous-for-submit/MAS_testing_repairing/tree/main/More_results/RQ2_all_result.pdf
From Table 4, we can find that removing any component degrades repair performance across all three MASs, indicating that all components contribute meaningfully and our design outperforms all alternative configurations.
Examining each component’s contribution reveals distinct patterns across MAS architectures. While disabling “Code Check” causes the smallest performance decrease, enabling “Code Check” still contributes to the overall repair performance and “Only Code Check” contributes approximately 40% of our method’s full repair capability, demonstrating that post-generation validation effectively catches misalignments even without enhanced plan interpretation. The relative importance of “Plan Interpretation” versus “Multi-prompt” varies by MAS architecture. In MetaGPT, “Only Plan Interpretation” substantially outperforms “Only Multi-prompt” because MetaGPT generates abstract plans that challenge coder comprehension, where interpretation bridges critical understanding gaps. Conversely, PairCoder benefits more from “Only Multi-prompt” since its plans have better readability, making the diversity of input expressions from multi-prompt generation more beneficial. These findings demonstrate that both strategies are valuable as different MASs exhibit unique flaws: MASs like MetaGPT with abstract plans benefit most from interpretation, while those like PairCoder with clear but rigidly expressed plans gains more from prompt diversity.
@@ 7.3. RQ3: Can our repairing method enhance the robustness of MASs against fuzzing?
From Fig. 5, we observe that on all three MASs, the speed of discovering failures (i.e., the slope of the blue line and the red line) slows down, and failure numbers decrease after applying our repairing method, showing that the repaired MASs exhibit superior robustness compared to their original counterparts. Notably, PairCoder (GPT-4o) exhibits the most remarkable robustness improvement against fuzzing, with 85.7% of failures in HumanEval ET eliminated. SCCG shows stable results across different backend LLMs and datasets, with 28.6%–53.8% decrease in failure numbers. These results indicate that our repairing method can effectively enhance the robustness of MASs against semantic-preserving mutations, resulting in fewer failures discovered during fuzzing.
Across different datasets, MASs exhibit the best robustness improvement on HumanEval ET, with failure numbers decreasing by 31.2%–85.7%, outperforming other datasets. Questions in HumanEval ET are well-defined and rarely contain complex logic, so our repairing method can effectively enhance the coder’s understanding of the plan by clearly identifying the question’s meaning, thereby improving MAS robustness.
Our repairing method enhances MAS robustness against fuzzing, with up to 85.7% failure number reductions. Repaired MASs exhibit superior robustness on the HumanEval ET dataset.
@@ 8. Discussion
Time Cost. Time efficiency is a critical factor in MASs, as computational overhead directly impacts system scalability and practical deployment (Chen et al., 2024b; Wu et al., 2023). To understand the time cost of our repairing method, we measure the execution time of MASs with and without our repairing method. Here, we select GPT-3.5 as the backend LLM and conduct experiments on HumanEval ET, as other settings (e.g., datasets, LLMs) show similar trends. Experiment results are shown in Table 6.
Table 6. Comparison of time cost on HumanEval ET with GPT 3.5 as backend LLM.
@@
In general, the additional time overhead primarily comes from the two components of our repairing method: multi-prompt generation and monitor agent insertion. For multi-prompt generation, since we maintain the same total number of generation attempts $n$, the primary overhead comes only from generating $k$ mutated prompts, which incurs little additional overhead for API calls compared to the main generation process. Besides, the monitor agent introduces two additional API calls per generation attempt, taking merely 2.7s–3.7s across different MASs.
We clarify that this overhead is modest in real-world application scenarios, given that a full MAS execution typically takes more than 30 seconds. On the other hand, the adoption of the monitor agent brings substantial robustness improvements, as shown in Sec. 7, making the cost-benefit trade-off favorable. In this regard, users of MASs are recommended to adopt our repairing method in most scenarios, especially for cases with complex requirements where robustness is their primary concern.
Threats to Validity. We identified and mitigated several potential threats to validity in our evaluation. The first threat is about the representativeness of our experimental setup. To address this, we evaluated three popular MASs (Dong et al., 2024; Zhang et al., 2024a; Hong et al., 2023) with three distinct LLMs (OpenAI, 2022; Achiam et al., 2023; Zhu et al., 2024) from different model families across four diverse datasets (Dong et al., 2025; Li et al., 2022; Yu et al., 2024b). The second threat is about the reliability of semantic-preserving mutation. To mitigate this threat, we employed GPT-4o (Achiam et al., 2023) to execute four carefully designed mutation operators. We conducted manual verification to confirm that the generated questions preserve the core semantics and thus the robustness issues revealed by these tests are meaningful. The last threat is subjectivity in manual failure analysis. To ensure reliable implications from failure cases, we first build a failure taxonomy based on a pilot study on 5% cases with the two software developers with over five years of programming experience. They then independently categorize the reasons for each failure for the other 15% cases, cross-check the annotation, and resolve all disagreements.
@@ 9. Related Work
MASs for Code Generation. Recent advancements in LLMs (Achiam et al., 2023; Zhu et al., 2024; Bai et al., 2023) have exhibited remarkable capabilities in code generation. By defining different roles and establishing communication protocols among multiple LLM agents, researchers have designed various MASs with diverse architectures (Hong et al., 2023; Dong et al., 2024; Zhang et al., 2024a; Huang et al., 2023; Islam et al., 2024; M. A. Islam, M. E. Ali, and M. R. Parvez, 2025). For instance, Hong et al. (Hong et al., 2023) propose MetaGPT, which employs five distinct agents to simulate the development workflow of a software company. Dong et al. (Dong et al., 2024) propose the Self-collaboration Code Generation framework, which includes three different agents responsible for planning, coding, and testing. To enhance the correctness of generated plans, some approaches (Zhang et al., 2024a; Islam et al., 2024; M. A. Islam, M. E. Ali, and M. R. Parvez, 2025) generate multiple candidate plans and then use a selection mechanism to identify the optimal plan with the highest score. For example, Zhang et al. (Zhang et al., 2024a) propose a clustering method to select the optimal plan, while Islam et al. (Islam et al., 2024) introduce a dynamic traversal method to assign the confidence of the generated plans. These approaches improve plan quality in terms of logical correctness, thereby outperforming previous approaches that rely on direct plan generation.
However, these methods focus primarily on logical correctness while neglecting the communication gap between planner and coder, which is first identified in this paper. In this work, we propose a repairing method to address the robustness issue of MASs by mitigating the gap.
Testing Code Generation Models. Numerous studies have applied testing methodologies to evaluate code generation LLMs. Wang et al. (Wang et al., 2023b) introduce ReCode, a benchmark for robustness evaluation through systematic perturbation testing, while Mastropaolo et al. (Mastropaolo et al., 2023) conduct empirical studies on GitHub Copilot’s robustness in real-world scenarios. Recent advances have proposed diverse methods to evaluate code generation models under different scenarios, like instruction concretization (Yan et al., 2025), multi-language robustness assessment (Rabbi et al., 2025), adversarial training (Zhang et al., 2024c), and equivalence checking (Wei et al., 2025). However, none of these works focus on MASs and their robustness remains under-explored. We conducted the first fuzzing-based testing approach to investigate the robustness of MASs and identified the planner-coder gap as the main cause of robustness issues.
@@ 10. Conclusion
This paper presents the first comprehensive study on the robustness of MASs for code generation, revealing that MASs fail to solve problems they initially solved when faced with semantically equivalent mutations. We identify the “planner-coder gap” as a fundamental cause of the robustness issues and propose a novel repairing method incorporating multi-prompt generation and monitor agent insertion. Our findings provide essential insights for developing more reliable MASs and establish a foundation for future research on robustness in multi-agent code generation systems.
11. Data Availability
We release our code and data to facilitate future research on MAS robustness at https://github.com/anonymous-for-submit/MAS_testing_repairing.

### v2

@@ 6.2. Overview of Our Repairing Method
As illustrated in Fig. 2, upon receiving an input requirement, we first mutate the requirement and generate multiple semantic-equivalent expressions to reduce likelihood of MASs misinterpreting specific expressions. For each generation, the planner generates the original plan and forwards it to the monitor, a new agent we introduce to mitigate the planner-coder gap. After receiving the plan from the planner, the monitor agent interprets the plan by providing detailed explanations for core concepts, edge cases, complex logic, relational phrases, and conditional judgments (see Sec. 5), which compensate for the information loss. Subsequently, the coder implements code based on this interpreted plan. In addition, the monitor also checks the implementation to enhance the alignment with the interpreted plan and requests regeneration when misalignments are detected.
@@ 6.4. Monitor Agent Insertion
Recent studies show that LLMs can provide evaluation highly aligned with humans (Wang et al., 2025; Zheng et al., 2023). Building upon this insight, we introduce a monitor agent between planner and coder to enhance plan quality and strengthen planner-coder alignment. This new agent performs two critical tasks: plan interpretation and code check, serving as a crucial mechanism to compensate for the information loss inherent in the code generation process. Note that this design is motivated by recent findings that specialized agents executing well-defined coordination and evaluation subtasks achieve superior reliability (Fourney et al., 2024; Valmeekam et al., 2023). By focusing on these well-defined, atomic tasks, the monitor exhibits greater robustness compared to the multi-stage transformations between planner and coder that trigger the planner-coder gap.
@@
To effectively address this issue and enhance plan quality, we prompt the inserted monitor agent to interpret the plan generated by the planner to make it more comprehensible to the coder. The monitor agent aims to enhance the clarity and completeness of the planned coding logic and mitigate the planner-coder gap, rather than correcting the semantic content of the plan. Specifically, the monitor interprets the decomposed requirement $\mathcal{R}$ and coding logic $\mathcal{L}$ with detailed explanations and clarifications targeting the five error patterns identified in Sec. 5, producing an interpreted plan $p^{\prime}$ that bridges the semantic gap between planning and coding. Fig. 3 presents an illustrative example of our prompt for the monitor agent. By interpreting the plan $p$ into $p^{\prime}$, the monitor facilitates the restoration of critical implementation details lost in Eq. 34 before code generation, which effectively reduces the information loss between the coding logic and final implementation. We employ few-shot prompting with carefully crafted examples to guide the interpretation process, focusing on the fundamental components of programming plans—such as core concepts, logic flows, and edge cases—rather than being tailored to any specific problem domain. This task-agnostic design provides general applicability across different code generation scenarios.
Clarification. Note that one might consider incorporating the monitor’s interpretation logic directly into the planner’s prompt. However, our proposed approach of a separate monitor agent offers advantages in terms of modularity and scalability. Rather than modifying the planner’s core functionality, this separation of concerns allows the monitor to be seamlessly integrated as a “quality gate” into various existing MAS architectures without altering their fundamental agent logic or task distribution. This design promotes reusability, easier maintenance, and addresses robustness issues across different MAS implementations without requiring architecture-specific modifications. We also follow the practice of consistent backend in a MAS to build the monitor agent with the backend LLMs of the MAS.
@@
Instead of using the testcase-based testing approaches that require dynamic execution with proper running environment, we directly prompt LLMs to perform static inspection and evaluate code compliance with the detailed specifications in $p^{\prime}$, which has been shown to be effective with the rich contextual information provided by the interpreted plan (Tong and Zhang, 2024; Zhuo, 2024). After the coder produces the solution code, we feed both the interpreted plan and the code back to the monitor to check whether the code complies with the plan regarding the different aspects of interpretation, as shown in Fig. 3(c). If the monitor identifies mismatches between code implementation and the plan, it returns the implementation to the coding agent for revision. By checking alignment between the interpreted plan $p^{\prime}$ and generated code $c$, the monitor creates a validation loop that detects and corrects information drift, effectively reducing the information loss that would otherwise accumulate through unchecked transformations. For efficiency, we use zero-shot prompting for code check and limit the process to execute only once.
@@ 7.1. RQ1: Can our repairing method effectively mitigate failures identified in fuzzing?
MAS Backend LLM HumanEval MBPP CodeContest CoderEval
Total Solved Rate Total Solved Rate Total Solved Rate Total Solved Rate
@@
Category EP-1 EP-2 EP-3 EP-4 EP-5 Overall Plan Logic Errors Invalid
Solved (%) 80.0 83.3 80.0 100.0 85.7 83.9 57.1 0.0
@@
To further understand the strengths of our method, Table 5 presents the repair results for different EPs identified in SCCG (GPT-3.5). Our repairing method demonstrates remarkable effectiveness in handling failures originating from the planner-coder gap, successfully resolving 83.9% of such cases. Notably, the consistently high repair ratios across all five EPs (80.0%–100.0%) indicate that our repairing method serves as a generalizable strategy for bridging the planner-coder gap, regardless of the specific type of error pattern involved. Surprisingly, our method also demonstrates considerable effectiveness on plan logic errors, successfully repairing 57.1% of such cases. Although our approach primarily targets planner-coder gap, this result suggests that the interpretation process may also compensate for the planner’s misunderstandings, as explicit interpretation may inadvertently correct or clarify flawed reasoning in the original plan.
@@ 7.2. RQ2: How do different components contribute to the overall performance?
We conduct ablation studies to investigate the effectiveness of multi-prompt generation and monitor agent insertion in our repairing method. For monitor agent insertion, we separately examine “Plan Interpretation” and “Code Check” to provide in-depth analysis. We present results with GPT-3.5 as the backend LLM, as it represents a typical configuration with other LLMs exhibiting similar patterns.
Fig. 4 shows that removing any component degrades repair performance across all three MASs, indicating that all components contribute meaningfully and our design outperforms all alternative configurations. While disabling “Code Check” causes the smallest performance decrease, enabling “Code Check” still contributes to the overall repair performance and “Only Code Check” contributes approximately 40% of our method’s full repair capability, demonstrating that post-generation validation effectively catches misalignments even without enhanced plan interpretation.
Moreover, our results reveal distinct patterns across MAS architectures, where the relative importance of “Plan Interpretation” versus “Multi-prompt” varies by MAS implementations. In MetaGPT, “Only Plan Interpretation” substantially outperforms “Only Multi-prompt” because MetaGPT generates abstract plans that challenge coder comprehension, where interpretation bridges critical understanding gaps. Conversely, PairCoder benefits more from “Only Multi-prompt” since its plans have better readability, making the diversity of input expressions from multi-prompt generation more beneficial. These findings demonstrate that both strategies are valuable as different MASs exhibit unique flaws: MASs like MetaGPT with abstract plans benefit most from interpretation, while those like PairCoder with clear but rigidly expressed plans gains more from prompt diversity.
@@ 7.3. RQ3: Can our repairing method enhance the robustness of MASs against fuzzing?
From Fig. 5, we observe that on all three MASs, the speed of discovering failures (i.e., the slope of the blue line and the red line) slows down, and identified failures decrease after applying our repairing method, showing that the repaired MASs exhibit superior robustness compared to their original counterparts. Notably, PairCoder (GPT-4o) exhibits the most remarkable robustness improvement against fuzzing, with 85.7% of failures in HumanEval eliminated. SCCG shows stable results across different backend LLMs and datasets, with 28.6%–53.8% decrease in failure numbers. These results indicate that our repairing method can effectively enhance the robustness of MASs against semantic-preserving mutations, resulting in fewer failures discovered during fuzzing.
Across different datasets, MASs demonstrate the most significant robustness improvement on HumanEval, with failures decreasing by 31.2%–85.7%, surpassing other datasets. This superior performance can be attributed to HumanEval’s well-defined questions with clear logic structures. Therefore, our repairing method effectively enhances the coder’s comprehension of the tasks by accurately capturing the question’s intent, thereby substantially improving MAS robustness.
Our repairing method enhances MAS robustness against fuzzing, with up to 85.7% failure number reductions. Repaired MASs exhibit superior robustness on the HumanEval dataset.
@@ 8. Discussion
Time Cost. Time efficiency is a critical factor in MASs, as computational overhead directly impacts system scalability and practical deployment (Wu et al., 2023). To understand the time cost of our repairing method, we measure the average execution time per attempt with and without our repairing method. Table 6 presents the experiment results on representative setting (HumanEval, GPT-3.5), as other settings (e.g., datasets, LLMs) show similar trends.
Table 6. Comparison of time cost on HumanEval with GPT 3.5 as backend LLM.
@@
In general, the additional time overhead primarily comes from the two components of our repairing method: multi-prompt generation and monitor agent insertion. For multi-prompt generation, since we maintain the same total number of generation attempts $n$, the primary overhead comes only from generating $k$ mutated prompts, which incurs little additional overhead for API calls compared to the main generation process. Besides, the monitor agent introduces two additional API calls per generation attempt, taking merely 2.7s–3.7s across different MASs. We clarify that this overhead is modest in real-world application scenarios, given that a full MAS execution typically takes more than 30 seconds. On the other hand, the adoption of the monitor agent brings substantial robustness improvements, as shown in Sec. 7, making the cost-benefit trade-off favorable.
Threats to Validity. We identified and mitigated several potential threats to validity in our evaluation. The first threat is the representativeness of our experimental setup. To address this, we evaluated three popular MASs (Dong et al., 2024; Zhang et al., 2024a; Hong et al., 2024) with three distinct LLMs (OpenAI, 2022; Achiam et al., 2023; Zhu et al., 2024) from different model families across four diverse datasets (Chen et al., 2021a; Austin et al., 2021; Li et al., 2022; Yu et al., 2024a). The second threat is the reliability of semantic-preserving mutation. To mitigate this threat, we conducted manual verification confirming that 99.2% of mutants preserve semantics. Additionally, we measure the semantic similarity using Sentence-Transformers (Reimers and Gurevych, 2019), whose results show 98.4% average similarity. Thus, the robustness issues revealed by our fuzzing are meaningful. The last threat is subjectivity in manual failure analysis. To ensure reliable implications from failure cases, two software developers with over five years of programming experience independently categorize failure reasons, cross-check the annotation, and resolve all disagreements.
@@ 9. Related Work
MASs for Code Generation.  Researchers have designed various MASs with diverse architectures by defining different roles among multiple LLM agents (Hong et al., 2024; Dong et al., 2024; Zhang et al., 2024a; Islam et al., 2024, 2025; Jiang et al., 2024b). For instance, MetaGPT (Hong et al., 2024) employs five distinct agents to simulate a software company’s development workflow, while the Self-Collaboration (Dong et al., 2024) uses three agents for planning, coding, and testing. To enhance plan correctness, Zhang et al. (Zhang et al., 2024a) use clustering to select the optimal plan, and Islam et al. (Islam et al., 2024) employ dynamic traversal to assign confidence scores to plans. However, these methods focus primarily on logical correctness while neglecting the communication gap between planner and coder, which is first identified in this paper. In this work, we propose a repairing method to address the robustness issue of MASs by mitigating the gap.
Testing Code Generation Models. Numerous studies have applied testing methodologies to evaluate code generation LLMs. Wang et al. (Wang et al., 2023a) introduce ReCode, a benchmark for robustness evaluation through systematic perturbation testing, while Mastropaolo et al. (Mastropaolo et al., 2023) conduct empirical studies on GitHub Copilot’s robustness in real-world scenarios. Recent advances have proposed diverse methods to evaluate code generation models under different scenarios (Chen et al., ; Yang et al., 2025; Zhong and Wang, 2024; Chen et al., 2023), like adversarial training (Zhang et al., 2024b) and instruction concretization (Yan et al., 2025). However, none of these works focus on MASs and their robustness remains under-explored. We conducted the testing-based study over MAS robustness and identified the planner-coder gap as the main cause of robustness issues.
@@ 10. Conclusion
This paper presents the first comprehensive study on the robustness of MASs for code generation, revealing their robustness issue on semantically equivalent mutations. We identify the “planner-coder gap” as the root cause and propose a novel repair method incorporating multi-prompt generation and monitor agent insertion. Our findings provide essential insights for developing more reliable MASs and establish a foundation for future robustness research in this domain.